OpenStudy (sloppycanada):

The length of the hypotenuse of a 45°-45°-90° triangle is 8 units. Find the exact value of the length of the leg.

Vocaloid (vocaloid):

leg = hypotenuse/(square root of 2)

OpenStudy (sloppycanada):

\[8\sqrt{2}\]

Vocaloid (vocaloid):

/ means divide

Vocaloid (vocaloid):

8 divided by sqrt(2)

OpenStudy (sloppycanada):

\[8\]

Vocaloid (vocaloid):



OpenStudy (sloppycanada):

\[\frac{ 8 }{ \sqrt{2} }\]

Vocaloid (vocaloid):

good, we can simplify that as 4*sqrt(2)

Vocaloid (vocaloid):